Find the general integral of the partial differential equation using Lagrange's method
p(z+e^x)+q(z+e^y)=z^2-e^(x+y)
1
Expert's answer
2020-11-24T08:08:19-0500

P=z+ex. Q=z+ey. R=z2 - ex+y

"\frac {dx} {z+e^x}= \frac {dy} {z+e^y} = \frac {dz}{z^2-e^{x+y}}"

Multiplyers: -z,e^x,1

"\frac {-zdx +e^xdy + dz} {-z^2 -ze^x + e^xz+ e^{x+y} + z^2-e^{x+y}} = \frac {-zdx +e^xdy + dz} {0}"

-zdx+exdy+dz=0

-zx+yex+z=c1


Multiplyers: ey,-z,1

"\frac {e^ydx -zdy + dz} {ze^y +e^{x+y}-z^2-ze^y+z^2-e^{x+y} } = \frac {e^ydx -zdy + dz} {0}"

eydx-zdy+dz=0

xey-zy+z=c2


Answer: c1=-zx+yex+z; c2=xey-zy+z